# DINO-Foresight: Looking into the Future with DINO

### v1

@@ 3.3 Compute-Efficient Training Strategies for High-Resolution Feature Forecasting
Using high-resolution features is crucial for pixel-wise scene understanding tasks, such as segmentation or depth prediction, where low-resolution features struggle to capture small objects or fine spatial structures [70, 17, 12, 81] To achieve good performance on these tasks, we aim to forecast VFM features extracted from frames with a spatial resolution of $H^{\prime}\times W^{\prime}=448\times 896$. For a ViT with a patch size of $14\times 14$, as used in DINOv2 [68] and EVA-CLIP [25], this results in feature maps with a resolution of $H\times W=32\times 64$, corresponding to 2048 tokens per frame.
However, training ViTs on such high-resolution inputs is computationally expensive in terms of both time and memory [23]. To address this challenge while maintaining efficient training, we explore the following strategies:
Low-Resolution Training with High-Resolution Inference
In this approach, we train on frames with a lower resolution of $224\times 448$, resulting in features with a resolution of $16\times 32$. During testing, we use high-resolution frames ($448\times 896$) and adapt the position embeddings through interpolation. However, this strategy leads to suboptimal performance due to a distribution shift between the training and test data, which causes inaccurate feature forecasting.
Sliding-Window Approach for High-Resolution Inference
Inspired by sliding-window techniques used in segmentation tasks [76], this strategy trains the model with cropped feature maps. The ViT encoder extracts features from high-resolution frames ($448\times 896$), producing high-resolution tokens (e.g., $32\times 64$ for a patch size of $14\times 14$). During training, we sample local crops of size $16\times 32$, taken from the same spatial locations across frames. The masked transformer is trained on these cropped features. During inference, the model applies the transformer in a sliding window fashion to the high-resolution features using these same local crops. This approach enables the model to efficiently handle large inputs without the computational burden of training on high-resolution features.
Two-Phase Training with Resolution Adaptation
Inspired by  [68, 81, 51], this strategy employs a two-phase training process. First, the model is trained on low-resolution frames ($224\times 448$) for several epochs, focusing on learning broad feature forecasting. Then, the model is fine-tuned on high-resolution frames ($448\times 896$) for a small number of epochs. This adaptation phase improves the model’s ability to handle high-resolution features without incurring the computational cost of training from scratch at the higher resolution.
As shown in our experiments, both strategies are effective, but the two-phase approach yields better feature forecasting performance. This is likely because the masked transformer has access to a larger spatial context when propagating VFM features in future frames.
@@ 3.4 Modular Framework for Future-Frame Multi-Task Predictions
Our framework is designed to work with a library of task-specific prediction heads, which can be easily applied to predicted future features after self-supervised training of the masked feature transformer. This approach is modular with respect to task heads: each head serves a different task and can be added or removed as needed without requiring retraining of the masked feature transformer.
We focus on three pixel-wise prediction tasks: semantic segmentation, depth prediction, and surface normals estimation. However, our framework can also support other scene understanding tasks, such as object detection and instance segmentation. For these tasks, we use the Dense Prediction Transformer [70] (DPT) architecture, which is well-suited to our setup. DPT is designed to work with multi-layer features from ViT-based encoders, aligning with the type of features that our masked feature transformer model learns to predict. DPT assembles these multi-layer ViT features, processes them at different resolutions, and progressively combines them into high-resolution predictions using convolutional layers. While we focus on DPT, our framework is flexible and can also accommodate other prediction heads, such as Mask2Former [17] or Mask R-CNN [36].
Prediction heads can be trained directly on frozen VFM features and then applied “off-the-shelf” to future-frame features predicted by the masked feature transformer. Additionally, they can be trained to account for the PCA stage by applying PCA compression and decompression to the multi-layer features. This approach is useful for cases where prediction heads are trained on annotated 2D images, without requiring video data, and then added to the library for future-frame predictions. Alternatively, prediction heads can be trained directly on the future-frame features predicted by an already pre-trained masked feature transformer, allowing them to adapt to potential distribution shifts between the original VFM features and the forecasted VFM features. In practice, we observed only minor performance differences between these two approaches.
@@ 4.1 Experimental setup
Data.  We assess our approach using the Cityscapes dataset [20], which offers video sequences of urban driving environments. The dataset includes 2,975 training sequences, 500 for validation, and 1,525 for testing, each with 30 frames captured at 16 fps and a resolution of 1024 × 2048 pixels. The 20th frame in each sequence is annotated for semantic segmentation with 19 classes. We train our model using the training sequences and evaluate it on the validation set.
@@
Implementation details.  By default, we use DINOv2-Reg with ViT-B/14 as the default VFM visual encoder for our method. For the masked feature transformer we built upon [6] implementation. We use 12 layers with a hidden dimension of $d=1152$ and sequence length $N=5$ (with $N_{c}=4$ context frames and $N_{p}=1$ future frame). For end-to-end training, we use the Adam optimizer [49] with momentum parameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate of $1.6\times 10^{-4}$ with cosine annealing. Training is conducted on 8 A100 40Gb GPUs with an effective batch size of 64. We train DPT heads for the semantic segmentation, depth prediction, and surface normals estimation tasks, and a Mask2Former head for the instance segmentation task. More implementation details Appendix C.
Evaluation Metrics.  To evaluate our method’s performance in predicting future semantic segmentation and depth maps, we use the following metrics: For semantic segmentation, we use mean Intersection over Union (mIoU) in two ways: (1) mIoU (ALL), which includes all semantic classes, and (2) MO-mIoU (MO), which considers only movable object classes like person, rider, car, truck, bus, train, motorcycle, and bicycle. For instance segmentation, we measure performance using average precision at a 0.50 IoU threshold (AP50) and also the mean average precision over IoU thresholds from 0.50 to 0.95. For depth prediction, we use two metrics: the mean Absolute Relative Error (AbsRel), defined as $\frac{1}{M}\sum_{i=1}^{M}\frac{|a_{i}-b_{i}|}{b_{i}}$, where $a_{i}$ and $b_{i}$ are the predicted and ground truth disparities at pixel $i$, and $M$ is the number of pixels. We also evaluate depth accuracy using $\delta_{1}$, the percentage of pixels where $\max\left(\frac{a_{i}}{b_{i}},\frac{b_{i}}{a_{i}}\right)<1.25$. For surface normal evaluation, we compute the mean angular error $\downarrow$ as $\frac{1}{N}\sum_{i=1}^{N}\cos^{-1}\left(\frac{\mathbf{n}_{i}\cdot\tilde{\mathbf{n}}_{i}}{\|\mathbf{n}_{i}\|\,\|\tilde{\mathbf{n}}_{i}\|}\right)$, where $\mathbf{n}_{i}$ and $\tilde{\mathbf{n}}_{i}$ are the predicted and ground truth normals, respectively. Furthermore, we measure precision through the percentage of pixels with angular errors below 11.25°, calculated as $\left(\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\theta_{i}<11.25^{\circ})\right)$.
Evaluation Scenarios  Following previous work [62, 66], we assess our model in two scenarios: short-term prediction (3 frames ahead,  0.18s) and mid-term prediction (9 frames ahead,  0.54s). In both cases, the target is the 20th frame. We subsample sequences by a factor of 3 before input to the model. For short-term prediction, the model uses frames 8, 11, 14, and 17 as context to predict frame 20 (with context length $N_{c}=4$ and $N_{p}=1$). For mid-term prediction, the model uses frames 2, 5, 8, and 11 as context and predicts frame 20 auto-regressively through frames 14 and 17. We calculate segmentation metrics on the 20th frame using Cityscapes ground truth. For depth and surface normals, we rely on pseudo-annotations from DepthAnythingV2 [94] and Lotus [35], respectively, due to the lack of true annotations in Cityscapes.
Method Short-term Mid-term
ALL MO ALL MO
3Dconv-F2F [18] 57.0 - 40.8 -
Dil10-S2S [62] 59.4 55.3 47.8 40.8
F2F [63] - 61.2 - 41.2
DeformF2F [73] 65.5 63.8 53.6 49.9
LSTM M2M [79] 67.1 65.1 51.5 46.3
APANet [43] - 64.9 - 51.4
F2MF [74] 69.6 67.7 57.9 54.6
PFA [58] 71.1 69.2 60.3 56.7
DINO Oracle 77.0 77.4 77.0 77.4
DINO-Foresight (ours) 71.8 71.7 59.8 57.6
Table 1: Comparison results with future semantic segmentation methods on Cityscapes validation set. ALL: mIoU of all classes. MO: mIoU of movable objects.
Method Short-term Mid-term
AP50 AP AP50 AP
Mask R-CNN Oracle 65.8 37.3 65.8 37.3
Mask H2F [63] 25.5 11.8 14.2 5.1
F2F [63] 39.9 19.4 19.4 7.7
CPConvLSTM [77] 44.3 22.1 25.6 11.2
APANet [43] 46.1 23.2 29.2 12.9
PFA [58] 48.7 24.9 30.5 14.8
DINO Oracle 56.1 32.1 56.1 32.1
DINO-Foresight (ours) 44.8 23.0 26.4 11.1
Table 2: Comparison with prior work on instance segmentation forecasting. DINO-Foresight and DINO Oracle use a Mask2Former prediction head.
Method Short-term Mid-term
ALL MO ALL MO
Segmenter [76] Oracle 78.6 80.8 78.6 80.6
Segmenter Copy-Last 55.5 52.7 40.5 32.2
VISTA + Segmenter 45.7 43.8 40.4 35.6
DINO Oracle 77.0 77.4 77.0 77.4
DINO Copy-last 54.7 52.0 40.4 32.3
DINO-Foresight (ours) 71.8 71.7 59.8 57.6
Table 3: Future prediction: VFM features vs RGB pixels. Results on semantic segmentation forecastig.
@@
Prediction 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
EVA2-CLIP [25] Oracle 71.0 69.5 71.0 69.5 85.2 .123 85.2 .123 3.37 94.5 3.37 94.5
@@
Prediction 65.3 59.3 52.5 43.9 81.3 .178 77.6 .209 3.80 92.8 4.49 89.2
@@
(c) 224$\rightarrow$224 N/A 64.50 62.63 55.49 52.62
(d) 224$\rightarrow$448 Pos. interp. 64.34 64.29 48.31 44.60
(e) 224$\rightarrow$448224 Sliding win. 71.26 71.11 58.75 56.78
(f) (224&\&&448$\rightarrow$448 Two-phase 71.81 71.71 59.78 57.65
4.2 VFM feature forecasting results
Comparison with Prior Work  In Tables 1 and 2, we compare our method to previous approaches in semantic and instance segmentation forecasting. On semantic segmentation forecasting, our method achieves superior results on all but one metric. For instance-segmentation forecasting, our method achieves performance comparable to the previous state-of-the-art, despite operating under a lower upper performance bound due to a weaker oracle (Mask R-CNN Oracle for prior work vs. DINO Oracle for us). This is likely due to the challenges of learning a strong instance segmentation model with a frozen visual encoder.
Unified Representations for Multiple Tasks: VFM Features vs. RGB Pixels
Our approach enables multiple scene understanding tasks with a single feature forecasting model, eliminating the need for task-specific models in previous methods, achieving competitive or superior results in semantic and instance segmentation forecasting (see Tables 1 and 2), as well as in depth prediction and surface normal estimation (see Table 4), underscoring its flexibility and practicality.
An alternative to forecasting VFM features is to predict future frames directly in RGB space, which also supports various downstream tasks. In this case, generated RGB frames can be processed with any scene understanding model. To compare these two approaches, we use VISTA [27], a recent large-scale video diffusion model for driving scenes. Given three context frames, VISTA generates 22 future frames. Using VISTA, we synthesize short-term and mid-term future frames and apply Segmenter [76] to these frames. Despite Segmenter’s superior oracle performance over our DPT head with frozen-DINOv2 features, VISTA achieves low semantic segmentation forecasting scores—even lower than Copy-Last (see Table 3). We provide an extended evaluation of our approach against the Copy-last and Vista baselines for depth and surface normals in Appendix Tab. 7. The Vista baseline only exceeds the Copy-last approach in the depth modality, while it produces poorer results in the surface normals modality.
Visual inspection (see Appendix Figs. 5 and 6) shows that VISTA achieves impressive quality and consistency for static parts of scenes or when there is relatively simple movement (e.g., all vehicles moving forward). However, it struggles with more complex motion, such as when vehicles move in different directions. This difficulty likely stems from VISTA’s need to model both low-level image details and high-level semantic changes, for which even its large-scale training proves insufficient (it was trained on 1740 hours of driving videos over 8 days on 128 A100 GPUs). In contrast, our method, which forecasts directly in the VFM feature space, abstracts away low-level pixel variations, enabling it to better forecast the semantic structure of the scene with far fewer training resources.
Comparison of VFM Visual Encoders  In Table 4, we evaluate our method using three distinct VFM encoders to extract the features on which our masked feature transformer operates (Prediction rows): DINOv2 with registers  [21, 68] (a self-supervised method), EVA2-CLIP [25] (a vision-language contrastive method), and SAM [50](a supervised instance segmentation method). For each, we use the ViT-B model variant. We also include results for the Copy-Last and Oracle baselines for comparison.
The results show that (1) DINOv2 consistently outperforms the other encoders across all tasks, providing the best results for both short-term and mid-term predictions. (2) This is in line with expectations, as the DINOv2-based Oracle also achieves the highest performance in all cases. (3) Additionally, our masked feature transformer proves effective in predicting future-frame features for all VFM encoders, showing significant improvement over the Copy-Last baseline. Based on these findings, we select DINOv2 as the default VFM visual encoder for our method.
@@ 4.2 VFM feature forecasting results
Surf. Normals Refer to caption Refer to caption Refer to caption Refer to caption
@@
Surf. Normals Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 4: Visualization of future predictions for semantic segmentation, depth, and surface normals. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
Training-Efficient Strategies for High-Resolution Feature Forecasting  In Table 5, we compare the resolution-adaptation strategies from Section 3.3, reporting results for future semantic segmentation. High-resolution features are essential: comparing the low-resolution Oracle baseline (model (a)) with the high-resolution Oracle baseline (model (b)) highlights the importance of high-resolution features for strong segmentation performance. Consequently, forecasting low-resolution features (model (c)) results in significantly poorer segmentation than models predicting high-resolution features (models (e) and (f)).
Adapting a model trained on low-resolution features for high-resolution inputs by simply adjusting position embeddings during inference (model (d)) leads to suboptimal results, even underperforming compared to low-resolution forecasting. The other two adaptation strategies—Sliding Window (model (e)) and two-phase training with resolution increase (model (f))—achieve considerably better results, demonstrating their effectiveness. The two-phase approach is simpler and yields the best performance, so we adopt it as our default for high-resolution feature forecasting.
4.3 Qualitative results.
4.4 Emerging Visual Representations in the Future-Frame Masked Feature Transformer
Self-supervised representation learning has achieved remarkable progress, with numerous studies focusing on extracting robust visual features from unlabeled images and videos [4, 80, 30, 72, 85, 13, 9, 37, 31, 88, 47, 46, 29]. Inspired by these advancements, we investigate the potential of our future-frame masked feature transformer as a self-supervised method for enhancing VFM visual features. Specifically, we train DPT heads for semantic segmentation and depth prediction, using not only the features predicted by the masked feature transformer but also additional features extracted from intermediate transformer layers. We examine features from the 6th, 9th, 10th, 11th, and 12th (last) layers of the transformer to assess whether these intermediate representations can further improve the strong VFM features predicted by our masked transformer.
Results, shown in Fig. 3, indicate that intermediate features from the transformer improve performance on both segmentation and depth forecasting tasks, except in one case (6th-layer features for short-term segmentation prediction). For segmentation, the best results are achieved with features from the 9th layer, while for depth, features from the 6th layer perform best. Although these improvements are modest, this aligns with expectations given the strength of the predicted VFM features alone.
While exploring self-supervised learning was not the primary aim of our work, we find these results intriguing, as they suggest that future prediction methods hold promise as self-supervised visual representation learners. We hope this work can spark further research into this direction.

### v2

@@ 3.3 Compute-Efficient Training Strategies for High-Resolution Feature Forecasting
Using high-resolution features is crucial for pixel-wise scene understanding tasks, such as segmentation or depth prediction, where low-resolution features struggle to capture small objects or fine spatial structures (Ranftl et al., 2021; Cheng et al., 2022; Chen et al., 2017; Touvron et al., 2019). To achieve good performance on these tasks, we aim to forecast VFM features extracted from frames with a spatial resolution of $H^{\prime}\times W^{\prime}=448\times 896$. For a ViT with a patch size of $14\times 14$, as used in DINOv2 (Oquab et al., 2024) and EVA-CLIP (Fang et al., 2024), this results in feature maps with a resolution of $H\times W=32\times 64$, corresponding to 2048 tokens per frame.
However, training ViTs on such high-resolution inputs is computationally expensive in terms of both time and memory (Dosovitskiy et al., 2020). To address this challenge while maintaining efficient training, we explore the following strategies:
Low-Resolution Training with High-Resolution Inference  In this approach, we train on frames with a lower resolution of $224\times 448$, resulting in features with a resolution of $16\times 32$. During testing, we use high-resolution frames ($448\times 896$) and adapt the position embeddings through interpolation. However, this strategy leads to suboptimal performance due to a distribution shift between the training and test data, which causes inaccurate feature forecasting.
Sliding-Window Approach for High-Resolution Inference  Inspired by sliding-window techniques used in segmentation tasks (Strudel et al., 2021), this strategy trains the model with cropped feature maps. The ViT encoder extracts features from high-resolution frames ($448\times 896$), producing high-resolution tokens (e.g., $32\times 64$ for a patch size of $14\times 14$). During training, we sample local crops of size $16\times 32$, taken from the same spatial locations across frames. The model is trained on these cropped features. At inference time, the model processes the high-resolution features in a sliding-window manner using the same crop dimensions. This approach efficiently handles large inputs while avoiding the computational cost of full-resolution training.
Two-Phase Training with Resolution Adaptation  This strategy employs a two-phase training process (Oquab et al., 2024; Touvron et al., 2019; Kolesnikov et al., 2020). First, the model is trained on low-resolution frames ($224\times 448$) for several epochs, focusing on learning broad feature forecasting. Then, the model is fine-tuned on high-resolution ($448\times 896$) for a small number of epochs. This adaptation phase improves the model’s ability to handle high-resolution features without incurring the computational cost of training from scratch at the higher resolution. As shown in our experiments, both strategies are effective, but the two-phase approach yields better feature forecasting performance. This is likely because the masked transformer has access to a larger spatial context when propagating VFM features in future frames.
@@ 3.4 Modular Framework for Future-Frame Multi-Task Predictions
Our framework supports a library of interchangeable task-specific prediction heads that operate on the preficted future freames, enabling flexible multi-task future understanding. The design is modular: each head operates independently, allowing tasks to be added or removed without retraining the core feature prediction model.
We focus on four pixel-wise prediction tasks: semantic segmentation, instance segmentation, depth prediction, and surface normals prediction. However, our framework is general and can support other scene understanding tasks, such as object detection and panoptic segmentation. For semantic segmentation, depth prediction, and surface normals prediction, we use the Dense Prediction Transformer (Ranftl et al., 2021) (DPT) architecture, which is well-suited to our setup. DPT leverages multi-layer features from ViT-based encoders—consistent with the features predicted by our model—and progressively refines them into high-resolution predictions using convolutional fusion. For the instance segmentation task, we use a Mask2Former (Cheng et al., 2022) head.
Prediction heads can be trained directly on frozen VFM features and then applied “off-the-shelf” to future-frame features predicted by the model. Additionally, they can be trained to account for the PCA stage by applying PCA compression and decompression to the multi-layer features. This approach is useful for cases where prediction heads are trained on annotated 2D images, without requiring video data, and then added to the library for future-frame predictions.
@@ 4.1 Experimental setup
Data.  We assess our approach using the Cityscapes (Cordts et al., 2016) and nuScenes (Caesar et al., 2020) datasets, both offering video sequences from urban driving environments. The Cityscapes dataset includes 2,975 training sequences, 500 for validation, each with 30 frames captured at 16 fps and a resolution of 1024 × 2048 pixels. The 20th frame in each sequence is annotated for semantic segmentation with 19 classes. The nuScenes dataset comprises of 700 training scenes and 150 validation scenes, captured at a frame rate of 12 Hz, with each sequence extending for 20 seconds.
@@
Implementation details.  By default, we use DINOv2-Reg with ViT-B/14 as the default VFM visual encoder for our method. For the masked feature transformer we built upon (Besnier and Chen, 2023) implementation. We use 12 layers with a hidden dimension of $d=1152$ and sequence length $N=5$ (with $N_{c}=4$ context frames and $N_{p}=1$ future frame). For end-to-end training, we use the Adam optimizer (Kingma and Ba, 2015) with momentum parameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate of $6.4\times 10^{-4}$ with cosine annealing. Training is conducted on 8 A100 40Gb GPUs with an effective batch size of 64. We train DPT (Ranftl et al., 2021) heads for the semantic segmentation, depth prediction, and surface normals estimation tasks, and a Mask2Former (Cheng et al., 2022) head for the instance segmentation task. More implementation details in Appendix Subsection D.
Table 1: Comparison with state-of-the-art on multiple forecasting tasks. Methods that do not handle a task are marked with ‘-’. For approaches requiring separate prediction models per task (e.g., PFA), we show multiple entries (semantic/instance). ALL: mIoU of all classes. MO: mIoU of movable objects. VISTAft is the VISTA model fine-tuned on Cityscapes. Compared approaches include: 3Dconv-F2F-RGB (Chiu et al., 2020), Dil10-S2S (Luc et al., 2017), F2F (Luc et al., 2018), ConvLSTM (Chiu et al., 2020), FeatReproj3D (Vora et al., 2018), Bayesian S2S (Bhattacharyya et al., 2019), 3Dconv-F2F-SEG (Chiu et al., 2020), DeformF2F (Šarić et al., 2019), LSTM AM S2S (Chen and Han, 2019), APANet (Hu et al., 2021), LSTM M2M (Terwilliger et al., 2019), IndRNN-Stack (Graber et al., 2021), DiffAttn-Fuse (Graber et al., 2022), F2MF (Saric et al., 2020), and PFA (Lin et al., 2021), Futurist (Karypidis et al., 2025) and VISTA (Gao et al., 2024).
Method Semantic Segmentation Instance Segmentation Depth Surface Normals
Short Mid Short Mid Short Mid Short Mid
ALL MO ALL MO AP50 AP AP50 AP $\delta_{1}$ $\delta_{1}$ 11.25 11.25
3Dconv-F2F-RGB 57.0 - 40.8 - - - - - - - - -
Dil10-S2S 59.4 55.3 47.8 40.8 - - - - - - - -
F2F - 61.2 - 41.2 39.9 19.4 19.4 7.7 - - - -
ConvLSTM 60.1 - - - - - - - - - - -
FeatReproj3D 61.5 - 45.4 - - - - - - - - -
Bayesian S2S 65.1 - 51.2 - - - - - - - - -
3Dconv-F2F-SEG 65.5 - 50.5 - - - - - - - - -
DeformF2F 65.5 63.8 53.6 49.9 - - - - - - - -
LSTM AM S2S 65.8 - 51.3 - - - - - - - - -
CPConvLSTM - - - - 44.3 22.1 25.6 11.2 - - - -
APANet - 64.9 - 51.4 46.1 23.2 29.2 12.9 - - - -
LSTM M2M 67.1 65.1 51.5 46.3 - - - - - - - -
IndRNN-Stack 67.6 60.8 58.1 52.1 - - - - - - - -
DiffAttn-Fuse 67.9 61.2 58.1 51.7 - - - - - - - -
F2MF 69.6 67.7 57.9 54.6 - - - - - - - -
PFA (semantic) 71.1 69.2 60.3 56.7 - - - - - - - -
PFA (instance) - - - - 48.7 24.9 30.5 14.8 - - - -
Futurist 73.9 74.9 62.7 61.2 - - - - 96.0 91.9 - -
Oracle 77.0 77.4 77.0 77.4 66.2 40.4 66.2 40.4 89.1 89.1 95.3 95.3
VISTAft 64.9 62.1 53.9 51.0 33.1 17.7 19.8 9.0 86.4 82.8 93.0 90.0
DINO-Foresight (ours) 71.8 71.7 59.8 57.6 50.5 26.6 27.3 12.6 88.6 85.4 94.4 91.3
[…]
Evaluation Scenarios  Following prior work (Luc et al., 2017; Nabavi et al., 2018), we evaluate our model on Cityscapes for short-term prediction (3 frames,  0.18s) and mid-term prediction (9 frames,  0.54s). On nuScenes, which has more static scenes (i.e., with less movement) than Cityscapes, we use mid-term prediction (9 frames,  0.75s) and long-term prediction (18 frames,  1.5s). Further details in Appendix Subsection F.
Baselines  We evaluate our method against three baselines. The first, the Oracle baseline, directly accesses the target future frame, establishing an upper performance bound. The second, Copy-Last, copies the most recent context frame to predict the target frame, providing a lower performance bound. Both baselines use DINOv2-Reg ViT-B encoder with DPT heads for semantic segmentation, depth prediction, and surface normals estimation, and a Mask2Former head for instance segmentation. The third baseline leverages VISTA (Gao et al., 2024), a state-of-the-art world model that uses video latent diffusion to generate future RGB frames from three context frames. This is a large-scale model, comprising 2.5 billion parameters and trained on 1,740 hours of driving videos, which we fine-tune on Cityscapes and nuScenes using the same frame-rate as our model. VISTA generates future RGB frames (with action-free conditioning), which are processed by a DINOv2-Reg encoder with DPT heads (identical to other baselines) for semantic segmentation, depth and surface normals prediction.
4.2 VFM feature forecasting results
Comparison with State-of-the-ArtTable 1 compares our method with state-of-the-art approaches for semantic/instance segmentation, depth estimation, and surface normal forecasting. The results highlight our key advantage: a single feature prediction model that achieves competitive or superior performance across multiple scene understanding tasks. In contrast, prior works either require separate prediction models per task (Luc et al., 2018) or handle at most two tasks simultaneously (Karypidis et al., 2025). This demonstrates the flexibility and practicality of our VFM feature forecasting approach.
Unified Representations for Multiple Tasks: VFM Features vs. RGB Pixels  An alternative to forecasting VFM features is predicting future frames directly in RGB space, which also supports performing multiple downstream tasks through standard scene understanding models. For comparison, we fine-tune VISTA (Gao et al., 2024) to generate five future frames from three context frames (8 frames in total). We process these synthesized frames (both short-term and mid-term) with DINOv2-Reg ViT-B and DPT heads (similar to our method’s setup) for fair evaluation. Despite VISTA’s large-scale training and model size (2.5B parameters), it achieves lower performance on semantic segmentation, depth estimation, and surface normal prediction. Extended evaluations on nuScenes (provided Appendix Table 7 ) show similar trends for depth and surface normal estimation. Our approach is also far more computationally efficient: mid-term forecasting on Cityscapes’ 500 validation scenes takes approximately 5 minutes versus VISTA’s 8.3 hours (both on a single A100 GPU). This highlights our method’s advantages of operating in VFM feature space – achieving accurate semantic prediction while being significantly more resource-efficient.
@@ 4.2 VFM feature forecasting results
Oquab et al. (2024) Prediction 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
EVA2-CLIP Oracle 71.0 69.5 71.0 69.5 85.2 .123 85.2 .123 3.37 94.5 3.37 94.5
@@
Kirillov et al. (2023) Prediction 65.3 59.3 52.5 43.9 81.3 .178 77.6 .209 3.80 92.8 4.49 89.2
VAE Oracle 47.3 34.7 47.4 35.2 61.5 .251 61.5 .251 5.3 86.1 5.3 86.1
Copy Last 37.1 25.6 28.5 16.5 60.7 .252 59.2 .286 5.8 83.2 6.3 80.1
Rombach et al. (2022) Prediction 33.4 17.9 24.7 9.8 64.1 .281 61.4 .394 6.5 80.5 8.0 73.2
Comparison of VFM Visual Encoders  In Table 2, we evaluate our method using three VFM encoders to extract features for our feature prediction model: DINOv2 with registers (Darcet et al., 2024; Oquab et al., 2024) (self-supervised), EVA2-CLIP (Fang et al., 2024) (vision-language contrastive), and SAM (Kirillov et al., 2023) (supervised instance segmentation). For each, we use the ViT-B variant. We also include Copy-Last and Oracle baselines for comparison. The results show that: (1) DINOv2 consistently outperforms other encoders across all tasks, achieving the best results for both short- and mid-term predictions. (2) This aligns with expectations, as the DINOv2-based Oracle also performs best in all cases. (3) Our model effectively predicts future-frame features for all VFMs, significantly improving over the Copy-Last baseline. Based on these findings, we select DINOv2 as our default VFM encoder.
Table 3: Continuous vs. Discrete VFM Representations. Comparison of continuous DINOv2 features (our approach) against 4M’s DINOv2 tokenizer with discrete codes. Unlike other tables and for fair comparison with the 4M tokenized variant, we use the DINOv2 ViT-B w/o Reg model and extract features from the last layer only. Results on semantic segmentation forecasting.
Method Continuous Discrete (4M tokenized)
Short Mid Short Mid
$uparrow ALL\uparrow MO\uparrow$ $uparrow ALL\uparrow MO\uparrow$
Oracle 72.9 74.0 72.9 74.0 70.2 71.4 70.2 71.4
Copy Last 54.7 51.9 40.5 32.2 53.7 51.0 40.0 31.6
Prediction 68.9 69.3 57.3 55.0 61.7 60.9 53.7 51.0
@@
(c) $\rightarrow224$ N/A 64.50 62.63 55.49 52.62
[…]
Future Prediction: VFM features vs VAE-based latents  Additionally, in Table 2, we evaluate using VAE latents (Rombach et al., 2022) (used in latent generative models) instead of VFM features, training DPT prediction heads on these latents. Results are significantly worse, as expected, since these latents lack high-level information, and even DPT oracles perform poorly. This highlights the advantage of forecasting VFM features over VAE latents.
Discrete vs. Continuous VFM Representations  Recent work in generative modeling has explored both discrete and continuous representations for image and video generation (Chang et al., 2022; Yu et al., 2023; Yan et al., 2021; Razavi et al., 2019; Li et al., 2024; Tschannen et al., 2024). Recent findings favor continuous representations (Li et al., 2024; Tschannen et al., 2024), showing that removing vector quantization can improve generation quality while retaining the benefits of sequence modeling. To further investigate this in the context of VFM feature forecasting, we employ 4M’s pretrained DINOv2 tokenizer (Bachmann et al., 2024), which encodes DINOv2 features (without register tokens) into discrete codes from a vocabulary of size 8192. We train a quantized variant of DINO-Foresight (ViT-B14, single-layer features) using a cross-entropy loss to predict these discrete codes, which are decoded back to DINOv2 features at inference time. Results are reported in Table 3. While the discretized variant achieves comparable oracle performance to the continuous VFM feature case, our continuous VFM feature forecasting approach yields superior future semantic prediction results. These findings suggest that preserving the rich, continuous representations from VFMs—without quantization—offers clear advantages for dense semantic forecasting tasks.
Training-Efficient Strategies for High-Resolution Feature Forecasting  In Table 4, we compare the resolution-adaptation strategies from Section 3.3, reporting results for future semantic segmentation. High-resolution features are essential: comparing the low-resolution Oracle baseline (model (a)) with the high-resolution Oracle baseline (model (b)) highlights the importance of high-resolution features for strong segmentation performance. Consequently, forecasting low-resolution features (model (c)) results in significantly poorer segmentation than models predicting high-resolution features (models (e) and (f)). Adapting a model trained on low-resolution features for high-resolution inputs by simply adjusting position embeddings during inference (model (d)) leads to suboptimal results, even underperforming compared to low-resolution forecasting. The other two adaptation strategies—Sliding Window (model (e)) and two-phase training with resolution increase (model (f))—achieve considerably better results, demonstrating their effectiveness. The two-phase approach is simpler and yields the best performance, so we adopt it as our default for high-resolution feature forecasting.
Additional Ablations and Analysis  We conduct comprehensive ablation studies in the appendix to further validate our approach. First, in Appendix Subsection A.3, we demonstrate strong zero-shot generalization by training on Cityscapes and evaluating on nuScenes without fine-tuning. The resulting performance is only slightly worse than models trained directly on nuScenes, while surpassing all baselines. Second, in Appendix Subsection A.5, we investigate the impact of different masking strategies (random vs. full masking) and loss functions (L1, MSE, SmoothL1, and SmoothL1 with cosine similarity). Third, in Appendix Subsection A.6, we analyze scalability across model sizes (Small: 115M, Base: 258M, Large: 460M parameters) and data scale (Cityscapes alone vs. Cityscapes+nuScenes combined), demonstrating consistent performance improvements with increased capacity and data diversity. These results highlight the promising robustness and scalability of our VFM feature forecasting approach.
4.3 Qualitative results.
@@
Normals Refer to caption Refer to caption Refer to caption Refer to caption
@@
Normals Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 3: Future predictions for semantic segmentation, depth, and surface normals. Noisy segmentations at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted features, as evident in the clear depth and surface normal predictions.